OpenStudy (anonymous):

solve he system of equation 2x-y=3 3y+3x=-9

OpenStudy (acxbox22):

multiply the top equation by 3

OpenStudy (anonymous):

i got 0 as my answer

OpenStudy (acxbox22):

okay so if you multiply thr top equation by 3 you get 6x-3y=9 and the second equation is just 3x+3y=-9 if you add the two equations you cancel out the x and you are left with 9x=0 solve and you get x=0 plug 0 for x in any equation 2(0)-y=3 -y=3 y=-3 so your solution is x=0 and y=-3

OpenStudy (anonymous):

thankss